Calculate the Hall coefficient of sodium based on free electron model. Sodium has bcc structure and the side of the cube is 4.28 Ã….


1
Expert's answer
2021-09-18T14:58:42-0400

The Hall coefficient is


"R_h=\frac{1}{ne}."

Also, we know that


"\frac{\Mu}{\rho}=a^3\frac{N_A}{n}\space\rightarrow\space n=\frac{\rho a^3N_A}{\Mu},\\\space\\
R_H=\frac{\Mu}{e\rho a^3N_A}=\\\space\\
=\frac{23·10^{-3}}{(1.602·10^{-19})(971)(4.28·10^{-10})^3(6.022·10^{23})}=\\\space\\
=3.13·10^{18}\text{ m}^3/\text{C}."